OpenStudy (anonymous):

Choose the product. -7p 3(4p 2 + 3p - 1) A:-28p5- 21p4+ 7p3 B:28p6+ 21p3- 7p C:21p3+ 8p2+ 3p4- 8p D:-21p5+ 8p4- 3p3

OpenStudy (anonymous):

Is this the full question?

OpenStudy (anonymous):

I'm guessing that the first line is the question and the rest are the answers?

OpenStudy (anonymous):

\(-7p^3(4p^2+3p-1)\) Is that what the questions looks like?

OpenStudy (anonymous):

yeah sorry

OpenStudy (anonymous):

The answer is C

OpenStudy (anonymous):

No problem. You can use the Equation button to help you write equations in the future. So, what is the question? Do you know how to multiply numbers with exponents?

OpenStudy (anonymous):

D

OpenStudy (anonymous):

The answer is D sorry

OpenStudy (anonymous):

Wait, a second

OpenStudy (anonymous):

nope plz explain

OpenStudy (anonymous):

Just give me a moment before a make a mistake let me check

OpenStudy (anonymous):

Oh the 3 was raised to the power

OpenStudy (anonymous):

Ok. You will distribute the \(-7p^3\). The first multiplication will be \(-7p^3 \times 4p^2\). Start by multiplying the coefficients.

OpenStudy (anonymous):

Answer is A by the way

OpenStudy (anonymous):

While you are explaining can you verify if I am correct, thanks

OpenStudy (anonymous):

Then, multiply the variable. \(p^3 = p \times p \times p\) and \(p^2 = p \times p\) so, \(p^3 \times p^2 = p^5\). In other words, when multiplying - add the exponents.

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

uhh can you show everthing you did to the equation i kinda slow

OpenStudy (anonymous):

OK. Starting from the beginning. Do you understand how to distribute?

OpenStudy (anonymous):

i think so- 28p^5+-21p+-7p

OpenStudy (anonymous):

Almost. Consider this: \(-7p^3(4p^2+3p-1)\) = \(-7p^3 \times 4p^2 + -7p^3 \times 3p -1 \times -7p^3\)

OpenStudy (anonymous):

Do the multiplication first. Do you understand how \(-7p^3 \times 4p@ = -28p^5\)?

OpenStudy (anonymous):

\(-7p^3 \times 4p^2 = -28p^5\)

OpenStudy (anonymous):

-28^4+-21p+7p^2?

OpenStudy (anonymous):

I'm not sure what you are doing to get that answer. I can guide you through this problem, but you will need to answer my questions so I can help you.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Do you understand the distribution? (-7p^3 \times 4p^2 + -7p^3 \times 3p -1 \times -7p^3\)

OpenStudy (anonymous):

\(-7p^3 \times 4p^2 + -7p^3 \times 3p -1 \times -7p^3\)

OpenStudy (anonymous):

im sorry the exponent 3's you were using looked like 2's

OpenStudy (anonymous):

\(\huge{-7p^3 \times 4p^2 + -7p^3 \times 3p -1 \times -7p^3}\) Is that more clear?

OpenStudy (anonymous):

-28p^3-21p^3 x 3p -1 x -7p^3 -49p^3 x3p-1 x -7p^3 good so far?

OpenStudy (anonymous):

The small type can be difficult to read on smaller screens. Let's start with the first multiplication: \(\huge{-7p^3 \times 4p^2}\)

OpenStudy (anonymous):

Yes. Now move on to the next multiplication.

OpenStudy (anonymous):

-21p^3

OpenStudy (anonymous):

\(\huge{-7p^3 \times 3p}\) remember that when an exponent is not written, it is understood to be 1.

OpenStudy (anonymous):

-21p

OpenStudy (anonymous):

What about the exponent?

OpenStudy (anonymous):

-21p^4

OpenStudy (anonymous):

Yes. And the last multiplication?

OpenStudy (anonymous):

7p^4?

OpenStudy (anonymous):

Why an exponent of 4?

OpenStudy (anonymous):

because -1 has an exponent of 1?

OpenStudy (anonymous):

That is true, but the -1 term has no variable. You only add the exponents when the base is the same.

OpenStudy (anonymous):

oh so u multiply them 7p^3

OpenStudy (anonymous):

Yes. Now string them all together.

OpenStudy (anonymous):

-28p5- 21p4+ 7p3, thank you that was'nt so hard after all

OpenStudy (anonymous):

Everything is easy after you learn how to do it. :-) Would you like to do another similar problem to make sure you understand it?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\huge{4x^2(3x^4+2x^3-6x^2+5x-2})\]

OpenStudy (anonymous):

12x^8+8x^6-24x^4+20x^2-8x^2

OpenStudy (anonymous):

20x^14-24x^4+20x^2-8x^2

OpenStudy (anonymous):

-4x^10+20x^2-8x^2

OpenStudy (anonymous):

Did you distribute? Then multiply each term individually? None should be combined.

OpenStudy (anonymous):

You were closest on your first attempt.

OpenStudy (anonymous):

idk my head hurts

OpenStudy (anonymous):

With your first answer, you were almost correct. You multiplied the exponents instead of adding them.

OpenStudy (anonymous):

\(\huge{4x^2 \times 3x^4 = (3\times 4)x^{2+4} = 12x^6}\)

OpenStudy (anonymous):

If the rules are difficult to remember, just expand it out: \(\huge{x^4 \times x^2 = (x \times x \times x \times x) \times (x\times x)}\) \(\huge{ = x\times x \times x \times x\times x\times x} = x^6\)

OpenStudy (anonymous):

12x^6 +8x^5-24^4+9x^3-8x2?

OpenStudy (anonymous):

Almost perfect. \(4 \times 5 = 20\) but your exponents are correct.

OpenStudy (anonymous):

x^23?

OpenStudy (anonymous):

It seems like you understand this now, do you want to do another shorter problem to make sure?

OpenStudy (anonymous):

nah im good but if i close this question can i still veiw this,and can you awnswer another question for me?

OpenStudy (anonymous):

Sure. and here is some interesting reading on the topic: http://www.uhv.edu/ac/mathsci/pdf/exponents.pdf

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Your Welcome.